OpenStudy (naveenbhatia1312):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (onyxxelite):

hello

OpenStudy (naveenbhatia1312):

hello

OpenStudy (welshfella):

do you know the formula for nth term of an arithmetic sequence? this is one and common difference ix 8

OpenStudy (naveenbhatia1312):

no clue

OpenStudy (welshfella):

notice we can write the first term as 8n - 10 where n = 0 the second one would be 8(1) + 10 = -2 and third = 8(2) - 10 = 6 the last term is 8(15) - 10 = 120 - 10 = 110

OpenStudy (bme89):

I believe it would be: \[\sum_{1}^{n}8n _{i}-18\]

OpenStudy (bme89):

except the bottom should be i=1 rather than 1

OpenStudy (welshfella):

yes

OpenStudy (naveenbhatia1312):

thats not one of the answers

OpenStudy (welshfella):

and n would be = 16 if it we take 110 to be the last term

OpenStudy (bme89):

What about \[\sum_{i=0}^{n}8n _{i}-10\]

OpenStudy (welshfella):

yes thats how i was going to write it both are correct

OpenStudy (welshfella):

agin maybe n = 15 because last term is 110

OpenStudy (welshfella):

Oh he's gone anyway!!

OpenStudy (bme89):

lol, oh well